The odds of event A happening are 1:3. The probability of event B happening is 40%. If events A and B are mutually exclusive events, then determine P(A or B).
Answer #1

P(Event A) = 1/(1 + 3) = 1/4 = 0.25

P(Event B) = 0.40

Hence,

P(A or B) = P(A) + P(B) - P(A and B) = 0.25 + 0.40 - 0 = 0.65